find the slope of the line whose y intercept is -9 and x intercept is 3
Answer by nerdybill(7384)

find the slope of the line whose y intercept is -9 and x intercept is 3
.
Point at y-intercept is (0, -9)
Point at x-intercept is (3, 0)
.
slope = (y2-y1)/(x2-x1)
slope = (0-(-9))/(3-0)
slope = (0+9)/(3-0)
slope = 9/3
slope = 3